# Retrieve to Explain: Evidence-driven Predictions with Language Models

### v2

@@ Appendix E Approximating Evidence Shapley Values
We used a simple Monte Carlo method to approximate Shapley values, combined with antithetical sampling for variance reduction (Mitchell et al., 2022). The Shapley value was approximated as
@@
where $\phi_{i}$ is the approximate Shapley value of feature $i$ (an encoded query/evidence pair), $M$ is the chosen number of sampled permutations, $S_{j}$ is the set of features preceding $i$ in the j-$th$ permutation sample, $g(S_{j})$ is the Reasoner output when only the features $S_{j}$ are unmasked, $g(S_{j}\cup\{i\})$ is the Reasoner output when feature $i$ is unmasked in addition to $S_{j}$, and $\bar{S}_{j}$ corresponds to the set of features preceding $i$ in the reverse of the j-$th$ permutation sample (equivalently, the set of features following $i$ in the j-$th$ permutation sample). The sum of the Shapley values over features plus the score when all features are $\mathrm{NULL}$ equates to the final score. Depending on the purpose, we use either the post-sigmoid output or the logit score for $g$. We use $M=100$ whenever Shapley values are computed as part of this paper.
@@
        $\phi_{p^{j}[i]}\leftarrow\frac{j-1}{j}\phi_{p^{j}[i]}+\frac{1}{j}(s_{i}-s_{i-1})$Cumulative average of marginals for feature $p^{j}[i]$ across permutations}
@@ Appendix F Relationship between Multinomial and Binary Objectives
R2E is trained to predict the probability that a given query-entity pair is “true”, i.e. that it came from a real occurrence in the literature and was not randomly generated. Given the labels $L\in\{0,1\}$, the query (masked sentence) variable $Q$, the named entity answer variable $A$, the Reasoner parameters $\theta$ and the fixed Retriever parameters $\psi$, the model is trained to predict
@@ Appendix G Post-hoc Frequency Bias Correction as Trading off Log Probability and Mutual Information
PMI is widely used in the NLP community to measure associations between keywords in a corpus, based on their marginal occurrence counts and joint co-occurrence counts (Jurafsky & Martin, 2019). Similarly, we find that it offers a straightforward means of correcting for class imbalance after training the model.
@@
which reflects a log probability estimate when $c=0$ and a pointwise mutual information estimate when $c=1$. We found that the best performance in terms of MRR on the Held-out Biomedical Literature validation dataset (see Appendix B), was achieved with a partial correction of $c=0.5$. We refer to this as R2E-cor, and refer to the case with $c=0.0$ as R2E-uncor.
@@
Appendix H Evaluation Dataset Statistics
The total sizes of all test/evaluation datasets are shown in Table S4.
Table S4: Evaluation dataset statistics
Dataset
Subset Count
@@ Appendix H Evaluation Dataset Statistics
Gene Description Facts
60839
@@
Appendix I Comparison of Models on a Non-Stratified Held-out Biomedical Literature Dataset
For completeness, we evaluated R2E on a dataset of 20,000 randomly-sampled publicly-available abstract sentences published from 2020 onwards, obtaining an imbalanced dataset without stratification by gene in Genes. As expected, the frequency-based baseline performs significantly better here relative to the stratified dataset in Table 1. Ranking metrics show similar performance for R2E, MCS and MLM (Table S5). In comparison, on the stratified dataset results in Table 1, R2E markedly outperforms baselines. This suggests that R2E obtains superior performance on the Held-out Biomedical Literature specifically after accounting for class imbalance (improved further by the bias correction procedure) and similar performance when class imbalance is not taken into account.
Table S5: Non-stratified Held-out Biomedical literature: R2E ranking metrics on a random subsplit (not stratified by gene in Genes) of query sentences published from 2020 onwards (20,000 queries), for an R2E model trained and retrieving from data prior to 2020.
@@ Appendix I Comparison of Models on a Non-Stratified Held-out Biomedical Literature Dataset
Appendix J Architecture Ablation Experiments
@@
A Held-out Biomedical Literature validation set was used for ablations experiments, consisting of one sentence per gene in Genes sampled from publicly-available abstract sentences published from 2020 onwards (as described in Appendices B & D), for an R2E model trained and retrieving from data prior to 2020. Ablation comparisons were made without applying any post-hoc frequency bias correction. Results are shown in Table S6.
We assessed R2E performance with an off-the-shelf biomedical transformer for the Retriever, in place of our task specific Retriever (Sections 3.2 & 3.3). We ablated the R2E Retriever by substituting in a PubMedBERT model initialised with its published weights (Gu et al., 2021). We created sentence embeddings by taking the mean over outputs corresponding to [MASK] tokens. This Retriever had a larger hidden size with 768 dimensional query and evidence embeddings. The R2E Reasoner was therefore linearly scaled to match this hidden size. We found that the default R2E had better ranking metrics than R2E with PubMedBERT substituted in for the Retriever.
We also assessed a simplification of the R2E Reasoner (Section 3.4.3) by substituting the convolutional query-evidence encoder with a parameter-free Hadamard product between the query embedding and each evidence embedding. The Hadamard product was chosen in order to incorporate an inductive bias towards the cosine similarity. Compared to R2E with the convolutional layers substituted for a Hadamard product, default R2E improved across all ranking metrics.
@@
Table S6: Architecture ablations: Ablated versions of R2E-Uncorrected on a validation dataset consisting of one sentence per gene in Genes sampled from sentences published from 2020 onwards, while training on and retrieving from data prior to 2020. Hadamard: substituting the convolution layers of the Reasoner with a Hadamard product. PubMedBERT: substituting the Retriever for the PubMedBERT model.
Metric R2E-Uncor Hadamard PubMedBERT
MRR 0.181 0.166 0.134
MR 3210 3260 3606
h@10 0.262 0.253 0.207
h@200 0.443 0.441 0.389
Appendix K Temporal Training
We tried a temporal training setup as an attempt to improve R2E’s ability to infer on future (2020 onwards) literature in the Held-out Biomedical Literature dataset from prior (pre-2020) retrieved literature. Specifically, we restricted retrieval during training to biomedical literature evidence sentences published prior to the query training sentence. This resulted in only small shifts in ranking metrics (Table S7).
Table S7: Temporal training: Held-out Biomedical Literature R2E-Corrected ranking metrics when restricting retrieval to biomedical literature published prior to the query sentence, during inference alone verses during training and inference.
Metric Query year R2E-Cor with pre-query retrieval
Inference-only Train & Inference
2020 0.233 0.220
MRR 2021 0.223 0.212
2022 0.219 0.207
2020 2489 2381
MR 2021 2544 2462
2022 2623 2549
2020 0.333 0.325
h@10 2021 0.324 0.315
2022 0.312 0.302
2020 0.521 0.525
h@200 2021 0.509 0.508
2022 0.496 0.493
@@
Appendix L Further Details on Creation of Gene Description Facts Dataset
@@
We downloaded UniProt FTP server data version 2023_01 and extracted descriptions of protein functions for our set of protein-coding gene entities (Genes) from UniProt (Universal Protein Resource), used under CC BY (4.0), (Consortium, 2022) (by pulling “text” from UniProt entities with type “function” in the “comment” field). Each entry is a human-written description of function, and consists of one or more sentences.
@@ 5. You can only skip a sentence if it contains no information about the
Appendix M Further Details on Creation of Explanation Annotations for Gene Description Facts Dataset
We constructed Evidence Annotations for the Gene Description Facts dataset by having GPT-4 annotate query relevance for all evidence across 50 randomly sampled query-entity pairs (positive examples) and the same 50 queries but with a randomly sampled alternative entity (negative examples), resulting in 6400 query-evidence pairs with a binary annotation.
@@
Appendix N Examples of Evidence Annotations for Gene Description Facts Dataset
@@
    False positive: R2E assigns a large, positive Shapley value to a piece of evidence that GPT-4 labels as irrelevant. R2E seems to consider the mention of magnesium concentration in relation to [MASK] sufficient to support the fact that ”[MASK] can use Mg (2+)”.
@@
Appendix O Further Details on Creation of Clinical Trial Outcomes Dataset
@@
From the remaining drugs, we extracted therapeutic hypotheses, as described by a combination of a drug’s protein targets and the diseases the drug had been tested against. We discretized therapeutic hypotheses using the PharmaProjects assigned MeSH (https://www.ncbi.nlm.nih.gov/mesh/) and Entrez (Maglott et al., 2005) ontology identifiers for the genes and diseases respectively. Nelson et al. 2015 and Minikel et al. 2023 investigate the transition between all trial phases. We use only a subset that focuses on the in vivo efficacy of therapeutic hypotheses. As such, we kept only the therapeutic hypotheses related to drugs tested at Phase II or III, or pre-Registration, Registration or Launched with regulatory approval. We kept only the therapeutic hypotheses where there were no drugs in active development and therefore whose clinical efficacy could be determined.
@@
Appendix P Genetics Baseline for the Clinical Trial Outcomes Dataset
Data for the genetics baseline was downloaded from the supplementary data of Minikel et al. 2023 (https://github.com/ericminikel/genetic_support/tree/sio/data) and reproduced using the methodology described in Minikel et al. 2023, briefly summarised below.
In the supplementary data, table assoc.tsv contains the full set of genetic associations that were templated into natural language in Section 4.5.2. These already-curated genetic associations were filtered further as per Minikel et al. 2023, removing all rows with a “source” of ‘OTG‘ and an “l2g_share” $<$ 0.5.
There exists ontological mismatch between sources of genetic evidence and diseases referenced in the PharmaProjects data. As such, the Clinical Trial Outcomes dataset is joined to the genetic association data by matching exactly on gene identity, and on a measure of MeSH-MeSH similarity for diseases.
The table sim.tsv.gz contains a full list of pairwise MeSH - MeSH similarities used in this joining of datasets. The similarity measure is a composite information criterion measure of similarity on the MeSH ontology tree; see Minikel et al. 2023 for details.
@@
Appendix Q Further Details on Creation of Evidence Annotations for Clinical Trial Outcomes Dataset
We constructed Evidence Annotations for the Clinical Trial Outcomes dataset by having GPT-4 annotate (as relevant or irrelevant) all evidence for 50 Clinical Trial Outcome therapeutic hypotheses associated with trial success, as well as 50 with trial failures, both randomly sampled, resulting in 6400 query-evidence pairs with a binary annotation.
@@
Appendix R Examples of Evidence Annotations for Clinical Trial Outcomes Dataset
@@
Appendix S Details on Cleaning MeSH Terms For Templating
@@
Appendix T Further Results on Predicting Clinical Trial Outcomes and Genetic Evidence
Table S8 shows that the best performing R2E model retrieving from literature alone, was R2E with bias correction and without temporal training (see Appendix K). As such, in Table 3, results without temporal training and with bias correction are shown.
Table S9 shows that the trends in AUROC are similar if we restrict the post-2005 therapeutic hypotheses as per Minikel et al. 2023. We restrict therapeutic hypotheses to only those diseases with ’genetic insight’. Diseases were only kept if there was at least one genetic association between a gene and a disease with a similarity of $>$ 0.7, leaving 4,226 therapeutic hypotheses.
@@
The genetics baseline has a widely published Relative Success (RS) of approximately 2 in predicting Clinical Trial success. This RS estimate is derived from a set of high confidence predictions (500 positive predictions across the 4,226 therapeutic hypotheses) from the genetics baseline (using a similarity threshold of $>$0.8 as per Minikel et al. 2023). R2E-cor predicting on literature obtained an RS equivalent to the genetics baseline (1.98; 95% CI (1.78, 2.21)), at a score threshold of 0.6, but with almost twice as many positive predictions (991 positive predictions across the 4,226 therapeutic hypotheses).
When calculating AUROC, the genetics baseline ranks every disease-gene pair in the Clinical Trial Outcomes data, based on the similarity of the trial disease to the nearest (most similar) genetic baseline disease, rather than binarising above and below a similarity of 0.8. The combination of low AUROC but high RS can be explained by the genetics baseline performing badly against the majority of therapeutic hypotheses for which there is no direct genetics evidence (i.e. the similarity of the nearest disease with genetic evidence is $<$ 0.8).
Table S10 shows that in the multi-modal context (with a corpus of sentences from the biomedical literature and from the genetics data), forcing retrieval of genetics evidence does not change the AUROC.
Figure S4 shows that there is substantial variation in performance across disease areas and modality. We highlight that the magnitude of difference in performance between R2E retrieving from genetics alone and R2E retrieving from literature, varies by disease area. This may indicate disease areas for which alternative modalities could be valuable to template into text.
Figure S5 shows the distribution of MeSH-MeSH ontological similarity, as calculated by Minikel et al. 2023, for query-evidence pairs with high Shapley scores (see Section 4.6) and where genetics evidence was annotated as relevant by GPT-4 (n=527) . Note that when calculating RS in Minikel et al. 2023, the threshold similarity for linking between therapeutic hypotheses and genetic association data is 0.8. Some examples of the “soft” semantic matching between the query and the genetics evidence for MeSH pairs with a similarity $<$ 0.2 are: (erythrocyte count and anemia), (eosinophilia and asthma), (astrocytoma and brain neoplasms). This highlights the shortcomings of a universal threshold based on ontological similarity metrics versus semantic matching and reasoning through natural language.
Table S8: Temporal training and Clinical Trial Outcomes: AUROC for R2E-Uncorrected and R2E-Corrected with and without enforcement of retrieval of literature evidence pre-query during training. Retrieval at inference time remains the same (i.e. pre-2005 evidence).
Model Train-time retrieval
Standard Pre-query
R2E-uncor 0.634 0.626
R2E-cor 0.639 0.628
@@
Table S9: Clinical Trial Outcomes on therapeutic hypotheses with genetic insight: AUROC for R2E retrieving from literature-alone, genetics-alone, or both; in comparison to baselines, when subsetting therapeutic hypotheses just to those where the disease has at least one genetic association in the genetics baseline.
Model Modality AUROC

### v4

@@ Appendix E Approximating Evidence Shapley Values
We used a simple Monte Carlo method to approximate Shapley values, combined with antithetical sampling for variance reduction Mitchell et al. (2022). The Shapley value was approximated as
@@
where $\phi_{i}$ is the approximate Shapley value of feature $i$ (an encoded query/evidence pair), $M$ is the chosen number of sampled permutations, $S_{j}$ is the set of features preceding $i$ in the j-$th$ permutation sample, $g(S_{j})$ is the Reasoner output when only the features $S_{j}$ are unmasked, $g(S_{j}\cup\{i\})$ is the Reasoner output when feature $i$ is unmasked in addition to $S_{j}$, and $\bar{S}_{j}$ corresponds to the set of features preceding $i$ in the reverse of the j-$th$ permutation sample (equivalently, the set of features following $i$ in the j-$th$ permutation sample). The sum of the Shapley values over features plus the score when all features are $\mathrm{NULL}$ equates to the final score. Depending on the purpose, we use either the post-sigmoid output or the logit score for $g$. We use $M=100$ whenever Shapley values are computed as part of this paper. See Appendix H for profiling of Shapley computation.
@@
         $\phi_{p^{j}[i]}\leftarrow\frac{j-1}{j}\phi_{p^{j}[i]}+\frac{1}{j}(s_{i}-s_{i-1})$
$\triangleright$ Cumulative average of marginals for feature $p^{j}[i]$ across permutations
@@ Appendix F Relationship between Multinomial and Binary Objectives
R2E is trained to predict the probability that a given query-entity pair is “true", i.e. that it came from a real occurrence in the literature and was not randomly generated. Given the labels $L\in\{0,1\}$, the query (masked sentence) variable $Q$, the named entity answer variable $A$, the Reasoner parameters $\theta$ and the fixed Retriever parameters $\psi$, the model is trained to predict
@@ Appendix G Post-hoc Frequency Bias Correction as Trading off Log Probability and Mutual Information
PMI is widely used in the NLP community to measure associations between keywords in a corpus, based on their marginal occurrence counts and joint co-occurrence counts Jurafsky and Martin (2019). Similarly, we find that it offers a straightforward means of correcting for class imbalance after training the model.
@@
which reflects a log probability estimate when $c=0$ and a pointwise mutual information estimate when $c=1$. We found that the best performance in terms of MRR on the Held-out Biomedical Literature validation dataset (Appendix B), was achieved with a partial correction of $c=0.5$. We refer to this as R2E-cor, and refer to the case with $c=0.0$ as R2E-uncor.
@@
Appendix H R2E Inference Speed
We profiled R2E for both prediction and explanation. We used CPUs only, though GPUs could be used to achieve additional speed-ups by reducing the time taken for the forward pass.
H.1 Prediction
For prediction on CPUs, the MLM baseline took $\sim$140ms over one query on one core, obtaining scores for all 19,176 genes via a single forward pass. By comparison, the non-negligible components of R2E inference time are:
  1. 1.
    The batched forward pass over 19,176 query-evidence pairs (one for each gene), through the Reasoner - $\sim$7.4s on one core, and scales linearly with cores
  2. 2.
    Vector searches over the 19,176 FAISS indices corresponding to each gene, for the Retriever - $\sim$27s on one core,  1.5s on 40 cores or <0.15s if one core per index
Since the evidence is split into separate retrieval indices for each of the potential answers, the top evidence from each can be found in parallel. Therefore, search can generally scale more efficiently than for a traditional single FAISS index. To optimise inference, the forward pass should be run in batches while the search results for each potential answer are returned from each corresponding FAISS index. As a result, the total time is then largely defined by the maximum time for the above two stages of batched forward pass and vector search, given the relevant parallelisation.
These results assume exact brute force vector search (IndexFlatIP search indices from FAISS Johnson et al. (2019)) with a complexity of $O(nd)$, where $n$ is the number of vectors in the given search index and $d$ is the dimensionality of each vector. While vector search was not a bottleneck in our setup, if inference speed were a concern as the retrieval corpus scales, there are many out-of-the-box options for more efficient approximate nearest neighbour search indices, including within FAISS. The R2E profiling results above also assume access to a machine with $\sim$300GB memory for the FAISS indices; fast inference is achieved on widely available resources.
H.2 Explanation
For inference time explanations, we compute Shapley values using the permutation-based method detailed in Appendix E, using $M=100$ permutations (200 with antithetical sampling). With 64 evidence sentences retrieved for a given query, this results in 12,800 evidence set variations required to compute all 64 Shapley values. Therefore, <10 forward passes are required, with a reasonable batch size. Given the small size of the Reasoner module (2 million parameters), generating an explanation takes $\sim$5 seconds using a single CPU only.
We also note that more efficient methods exist for approximating Shapley values Lundberg and Lee (2017), particularly for deep networks. However, since Shapley value efficiency is neither our primary focus nor prohibitive, we used a permutation-based approach (Appendix E).
Appendix I Evaluation Dataset Statistics
The total sizes of all test/evaluation datasets are shown in Table 4.
Table 4: Evaluation dataset statistics
Dataset Subset Count
@@ Appendix I Evaluation Dataset Statistics
Gene Description Facts 60839
@@
Appendix J Predicting Genes in Held-out Biomedical Literature
Dataset Construction
For all experiments in this section, we trained the MLM (R2E Retriever) and R2E Reasoner only on biomedical literature data published prior to 2020. Except where specified, R2E also only retrieved data published prior to 2020 (Figure 3). We then constructed a Held-out Biomedical Literature evaluation dataset from publicly-available paper abstracts. We generated a balanced dataset to obtain results reflecting performance across all 19,176 genes, not biased to the most well-studied (discussed further in Appendix K). We sampled one sentence per unique gene in Genes for each of the years 2020, 2021, and 2022; further details in Appendix B.
Results
R2E improved on the baselines over all year subsets, both with and without bias correction (Table 5). Bias-corrected R2E improved on uncorrected performance, consistent with the use of a balanced evaluation dataset. For completeness, we show results on an imbalanced dataset (without stratification by gene in Genes) in Appendix K.
To test R2E’s ability to leverage retrieved literature that it was not trained on, we enabled retrieval up to the year preceding the query sentence publication (rather than strictly prior to the 2020 training data cutoff). This improved performance (R2E-cor-updated, Table 5).
Table 5: Held-out Biomedical Literature: Ranking metrics on a dataset consisting of one sentence per gene in Genes for each year of 2020, 2021, and 2022. MLM and R2E trained on data published prior to 2020. MCS, R2E-uncor and R2E-cor also retrieved data published prior to 2020. R2E-cor-updated retrieved up to the year before the publication year of the query sentence.
Metric Query year Baselines R2E
Freq MCS MLM Uncor Cor Cor-Updated
2020 $<$0.001 0.182 0.181 0.198 0.233 -
MRR 2021 $<$0.001 0.172 0.169 0.187 0.215 0.223
2022 $<$0.001 0.167 0.164 0.178 0.205 0.219
2020 7661 3280 3465 2803 2489 -
MR 2021 7834 3568 3789 3032 2695 2544
2022 7931 3770 4016 3287 2902 2623
2020 $<$0.001 0.268 0.269 0.291 0.333 -
h@10 2021 $<$0.001 0.251 0.252 0.274 0.313 0.324
2022 $<$0.001 0.243 0.243 0.260 0.295 0.312
2020 0.014 0.443 0.438 0.484 0.521 -
h@200 2021 0.014 0.422 0.416 0.456 0.497 0.509
2022 0.013 0.404 0.398 0.435 0.473 0.496
Appendix K Comparison of Models on a Non-Stratified Held-out Biomedical Literature Dataset
Gene mention counts are extremely imbalanced in the literature. In the training data, of the 19,176 protein-coding genes, the most-well studied has approximately 2 million mentions, while the least studied 10,000 genes all have less than 1,000 mentions. For our Held-out Biomedical Literature dataset we used stratified sampling (stratification by gene in Genes) to obtain a class balanced test dataset, with equal counts of each gene to avoid dominance of well-studied genes. By preventing reliance of models on the gene frequency distribution prior, a class-balanced setup is especially challenging. Strong performance across the genome is desirable because understudied genes are of particular interest in drug discovery, when seeking new ways to treat a disease.
While our focus is therefore on balanced performance across the genome (results in Appendix J), for completeness, we also evaluated R2E on a dataset of 20,000 randomly-sampled publicly-available abstract sentences published from 2020 onwards, obtaining an imbalanced dataset without stratification by gene in Genes. As expected, the frequency-based baseline performs significantly better here relative to the stratified dataset in Table 5, reflecting that ability to rely on the frequency distribution prior. Ranking metrics show similar performance for R2E, MCS and MLM (Table 6). In comparison, on the more challenging stratified setup R2E markedly outperforms baselines (Table 5). Comparing R2E and MLM, R2E’s superior balanced performance across the genome could be explained by it’s access to a knowledge base even for the most rare genes, avoiding the need to memorise knowledge of genes rarely seen at training time in the model parameters. R2E obtains superior performance on less studied genes without sacrificing performance on well-studied genes.
Table 6: Non-stratified Held-out Biomedical literature: R2E ranking metrics on a random subsplit (not stratified by gene in Genes) of query sentences published from 2020 onwards (20,000 queries), for an R2E model trained and retrieving from data prior to 2020.
@@ Results
Appendix L Architecture Ablation Experiments
@@
We performed ablations of all core R2E architectural components, including the Reasoner, Retriever and frequency bias correction. A Held-out Biomedical Literature validation set was used for ablations experiments, consisting of one sentence per gene in Genes sampled from publicly-available abstract sentences published from 2020 onwards (as described in Appendices B & D), for an R2E model trained and retrieving from data prior to 2020. The results are summarised in Table 7. All ablations resulted in a drop in performance across all ranking metrics, demonstrating the benefit of R2E components.
@@
Table 7: Architecture ablations: Ablated versions of R2E-uncor on a validation dataset consisting of one sentence per gene in Genes sampled from sentences published from 2020 onwards, while training on and retrieving from data prior to 2020. Hadamard: substituting the convolution layers of the Reasoner with a Hadamard product. PubMedBERT: substituting the Retriever for the PubMedBERT model.
Metric R2E Reasoner Ablations Retrieval Ablations
Cor Uncor MCS Hadamard PubMedBERT MLM
MRR 0.211 0.181 0.163 0.166 0.134 0.163
MR 2873 3210 3726 3260 3606 3945
h@10 0.302 0.262 0.241 0.253 0.207 0.242
h@200 0.482 0.443 0.409 0.441 0.389 0.404
L.1 Reasoner
The MCS baseline (Section 4.2) acts as an ablation of the R2E Reasoner, since it relies solely on query-evidence cosine similarities of the Retriever to obtain a score.
We also selectively ablated the convolutional query-evidence encoder component of the R2E Reasoner (Section 3.4) by substituting that component for a parameter-free Hadamard product between the query embedding and each evidence embedding. The Hadamard product was chosen in order to incorporate an inductive bias towards the cosine similarity.
L.2 Retriever
We ablated our task specific Retriever (Sections 3.2 & 3.3), by replacing it with an off-the-shelf biomedical transformer. Specifically we used a PubMedBERT model initialised with its published weights Gu et al. (2021) as the Retriever. We created sentence embeddings by taking the mean over outputs corresponding to [MASK] tokens. This Retriever had a larger hidden size with 768 dimensional query and evidence embeddings. The R2E Reasoner was therefore linearly scaled to match this hidden size.
We also evaluated the MLM baseline (Section 3.2), which acts as an ablation of R2E in its entirety, taking a fully parametric approach to prediction.
L.3 Post-hoc frequency bias correction
We report results with and without bias correction.
@@
Appendix M Further Details on Creation of Gene Description Facts Dataset
@@
We downloaded UniProt FTP server data version 2023_01 and extracted descriptions of protein functions for our set of protein-coding gene entities (Genes) from UniProt (Universal Protein Resource), used under CC BY (4.0), Consortium (2022) (by pulling “text” from UniProt entities with type “function” in the “comment” field). Each entry is a human-written description of function, and consists of one or more sentences.
@@ 5. You can only skip a sentence if it contains no information about the
Appendix N Further Details on Creation of Explanation Annotations for Gene Description Facts Dataset
We constructed Evidence Annotations for the Gene Description Facts dataset by having GPT-4 annotate query relevance for all evidence across 50 randomly sampled query-entity pairs (positive examples) and the same 50 queries but with a randomly sampled alternative entity (negative examples), resulting in 6400 query-evidence pairs (100 queries each retrieving 64 pieces of evidence) with a binary annotation. The same instructions were followed by the drug discovery expert when providing annotations used to validate the GPT-4 annotations in Section 4.4. The expert annotator was a Principal Scientist with over two years industry drug target identification experience in addition to holding a domain-relevant PhD and post-doc. They consented to the use of their annotations.
@@
Appendix O Examples of Evidence Annotations for Gene Description Facts Dataset
@@
    False positive: R2E assigns a large, positive Shapley value to a piece of evidence that GPT-4 labels as irrelevant. R2E seems to consider the mention of magnesium concentration in relation to [MASK] sufficient to support the fact that "[MASK] can use Mg (2+)".
@@
Appendix P Further Details on Creation of Clinical Trial Outcomes Dataset
@@
From the remaining drugs, we extracted therapeutic hypotheses, as described by a combination of a drug’s protein targets and the diseases the drug had been tested against. We discretized therapeutic hypotheses using the PharmaProjects assigned MeSH (https://www.ncbi.nlm.nih.gov/mesh/) and Entrez Maglott et al. (2005) ontology identifiers for the genes and diseases respectively. Nelson et al. 2015 and Minikel et al. 2024 investigate the transition between all trial phases. We use only a subset that focuses on the in vivo efficacy of therapeutic hypotheses. As such, we kept only the therapeutic hypotheses related to drugs tested at Phase II or III, or pre-Registration, Registration or Launched with regulatory approval. We kept only the therapeutic hypotheses where there were no drugs in active development and therefore whose clinical efficacy could be determined.
@@
Appendix Q Genetics Baseline for the Clinical Trial Outcomes Dataset
Data for the genetics baseline was downloaded from the supplementary data of Minikel et al. 2024 (https://github.com/ericminikel/genetic_support/tree/sio/data) and reproduced using the methodology described in Minikel et al. 2024, briefly summarised below.
In the supplementary data, table assoc.tsv contains the full set of genetic associations that were templated into natural language in Section 4.5. These already-curated genetic associations were filtered further as per Minikel et al. 2024, removing all rows with a “source” of ‘OTG‘ and an “l2g_share” $<$ 0.5.
There exists ontological mismatch between sources of genetic evidence and diseases referenced in the PharmaProjects data. As such, the Clinical Trial Outcomes dataset is joined to the genetic association data by matching exactly on gene identity, and on a measure of MeSH-MeSH similarity for diseases / traits.
The table sim.tsv.gz contains a full list of pairwise MeSH - MeSH similarities used in this joining of datasets. The similarity measure is a composite information criterion measure of similarity on the MeSH ontology tree; see Minikel et al. 2024 for details.
@@
Appendix R Further Details on Creation of Evidence Annotations for Clinical Trial Outcomes Dataset
We constructed Evidence Annotations for the Clinical Trial Outcomes dataset by having GPT-4 annotate (as relevant or irrelevant) all evidence for 50 Clinical Trial Outcome therapeutic hypotheses associated with trial success, as well as 50 with trial failures, both randomly sampled, resulting in 6400 query-evidence pairs (100 queries each retrieving 64 pieces of evidence) with a binary annotation. The same instructions were followed by the drug discovery expert when providing annotations used to validate the GPT-4 annotations in Section 4.5. The expert annotator was a Principal Scientist with over two years industry drug target identification experience in addition to holding a domain-relevant PhD and post-doc. The expert consented to the use of their annotations.
@@
Appendix S Examples of Evidence Annotations for Clinical Trial Outcomes Dataset
@@
Appendix T Details on Cleaning MeSH Terms for Templating
@@
Appendix U Further Results on Predicting Clinical Trial Outcomes and Genetic Evidence
U.1 Relative Success
For a given prediction threshold, we compute relative success of model predictions as:
$RS=\frac{(\text{True Positive}/\text{Predicted Positive})}{(\text{False Negative}/\text{Predicted Negative})}$ (16)
Where relevant, we use Katz method Katz et al. (1978) for confidence intervals and Z-test for comparisons.
U.2 Results for Diseases with Genetic Insight
Previous analyses of genetic methods for target identification have restricted to evaluating only on diseases with at least one piece of genetics data and for which therefore genetics could be expected to be informative (those with ’genetic insight’) Minikel et al. (2024). In Minikel et al. (2024), diseases were deemed to have genetic insight if there was at least one genetic association between a gene and disease with a MeSH-MeSH similarity of $>$ 0.7. This subsetting of therapeutic hypotheses was used to obtain the widely published relative success of $\sim$2 in predicting clinical trial outcome success from genetic data.
@@
We validated our Clinical Trial Outcomes dataset by corroborating this result by similarly restricting post-2005 therapeutic hypotheses to diseases with genetic insight, and using a MeSH-MeSH similarity threshold of $>$0.8 as the threshold for positive predictions as per Minikel et al. 2024. At this threshold, the genetics baseline makes 500 positive predictions across the 4,056 therapeutic hypotheses, with a Relative Success of 1.98, 95% CI (1.76, 2.24). In comparison, R2E-cor predicting on literature obtained a relative success of 2.17 (95 % CI (2.44, 1.93)) making the same number of positive predictions.
For completeness, we also show AUROC results after restricting to diseases with genetic insight in Table 8, with trends in AUROC similar to the results without restriction shown in the main text - rationale for the latter below (Appendix U.3).
@@
Table 8: Clinical Trial Outcomes on therapeutic hypotheses with genetic insight: AUROC for R2E retrieving from literature-alone, genetics-alone, or both; in comparison to baselines, when subsetting therapeutic hypotheses just to those where the disease has at least one genetic association in the genetics baseline.
Model Corpus AUROC